# Achieving Optimal Sample Complexity for a Broader Class of Signals in Sparse Phase Retrieval

### v2

@@ Appendix B Proof for Proposition 1
Denote $\overline{S}$ as the index set corresponding to the largest $p$ elements of $\mathbf{x}$. We establish the upper and lower bounds for $\frac{1}{p}\sum_{j\in S^{0}}Z_{j}$ and $\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}$, respectively.
$\frac{1}{p}\sum_{j\in S^{0}}Z_{j}\leq$ $~\frac{1}{p}\sum_{j\in S^{0}}\mathbb{E}\left[\tilde{Z}_{j}\right]+\left\lvert\frac{1}{p}\sum_{j\in S^{0}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
$\leq$ $~\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{4p\|\mathbf{x}\|^{2}}+\frac{1}{16ps(p)},$ (B.1)
$\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}\geq$ $~\frac{1}{p}\sum_{j\in\overline{S}}\mathbb{E}\left[\tilde{Z}_{j}\right]-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
$=$ $\frac{\|\mathbf{x}_{\overline{S}}\|^{2}}{4p\|\mathbf{x}\|^{2}}-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
@@
Thus we complete the proof.
@@ Appendix C Proof for Proposition 2
For any set $\mathcal{T}$} satisfying $|{\mathcal{T}}|=p^{\prime}\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$, suppose $\mathbf{z}$ as the unit eigenvector corresponding to the largest eigenvalue, denoted as $\tau$, of
@@
where $(a)$ is based on the relationship between matrix norm and vector norm, $(b)$ employs (A.9) with $g(k,s(p))=s(p)$ and the condition $\|\mathbf{z}\|=1$. The above inequality holds with probability at least $1-2e^{-c_{\eta}m}$ when $m\geq C_{\eta}p^{\prime}s^{2}(p)\log n$. Therefore, through simple transformation, (C.6) can be expressed as
@@
Moreover, since $\tau$ is the largest eigenvalue of $\mathbf{Z}_{{\mathcal{T}}}$, we can estimate it as
@@
where $(a)$ comes from the definition of matrix norm and Lemma 4 ((A.9)) with g(k,s(p))=s(p)g(k,s(p))=s(p) [^5] In fact, since the matrix $\mathbf{Z}_{TT}-\frac{\mathbf{x}_{{\mathcal{T}}}\mathbf{x}^{*}_{{\mathcal{T}}}}{4\|\mathbf{x}\|^{2}}$ are Hermitian matrix, all of its eigenvalues are real and non-negative, allowing the inequality hold true. . Taking (C.16) into (C.7) yields
@@
where $(a)$ comes from the assumption $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$. Let $\mu_{1}=\sqrt{(1-16\eta)}$, we complete the proof.
@@ Appendix D Proof for Proposition 3
$\overset{(a)}{\geq}$ $~\frac{1}{4\|\mathbf{x}\|^{2}}\times\frac{\gamma\|\mathbf{x}\|}{2\sqrt{k}}\times\mu_{1}\sqrt{\frac{1}{2s(p)}}\|\mathbf{x}\|-|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$ (D.7)
$\overset{(b)}{>}$ $~\frac{\gamma\mu_{1}}{8\sqrt{2}}\sqrt{\frac{1}{ks(p)}}-\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}}=\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.10)
where $(a)$ is from the definition of $S_{\gamma}$ and Proposition 2 and $(b)$ employs Lemma 6 with $\eta=\frac{\mu_{1}}{8\sqrt{2}}$ and holds with probability exceeding $1-4e^{-c_{\gamma,\mu_{1}}m}$ when $m\geq C_{\mu_{1}}\frac{ks(p)}{\gamma^{2}}\log n$.
@@
$|f_{l}|\leq|\mathbb{E}[\tilde{f}_{l}]|+|f_{l}+\mathbb{E}[\tilde{f}_{l}]|\overset{(a)}{\leq}\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.11)
where $(a)$ comes from (D.1) and Lemma 6 with $\eta=\frac{\mu_{1}}{8\sqrt{2}}$. By combining (D.10) and (D.11), we claim that all indices in $S_{\gamma}$ are selected in this case. Therefore, we can derive that
$\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\sum_{i\notin S_{\gamma}}\frac{|x_{i}|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}.$
@@ Appendix E Proof of Proposition 4
Let $\mathbf{z}$ with $\|\mathbf{z}\|=\lambda$ be the eigenvector corresponding to the largest eigenvalue of $\mathbf{Z}_{{\mathcal{T}}}$. Following the same analysis as in Appendix C with the set $\mathcal{T}$} satisfying $|{\mathcal{T}}|\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}$, we obtain similar conclusion to (C.7) and (C.16) as
$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}$ $\geq 4(\tau-\eta_{2})\|\mathbf{z}\|^{2}\|\mathbf{x}\|^{2},$ (E.1)
$\tau$ $\geq\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}}-\eta_{2},$ (E.2)
where we employ $\|\mathbf{z}\|=\lambda$ and Lemma 4 with $g(k,s(p))=1$. Hence (E.1) and (E.2) hold with probability at least $1-2e^{-c_{\gamma,\eta_{2}}m}$ if $m\geq C_{\gamma,\eta_{2}}k\log n$. Suppose $\kappa_{0}=\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}$, taking (E.2) into (E.1) yields
$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}\overset{(a)}{\geq}\left(\kappa_{0}-8\eta_{2}\right)(1-\eta_{1})\|\mathbf{x}\|^{4}\overset{\triangle}{=}\mu_{2}^{2}\|\mathbf{x}\|^{4},$ (E.3)
where $(a)$ employs Lemma 3 with $\kappa=1$ and $\beta=\eta_{1}$. In the following, We derive the estimation for $\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})$. From the property of inner product, we have
$\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}=$ $~\|\mathbf{x}_{{\mathcal{T}}}\|^{2}+\|\mathbf{z}\|^{2}-2|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|$
$\leq$ $~(1+\eta_{1}+\kappa_{0})\|\mathbf{x}\|^{2}-2\mu_{2}\|\mathbf{x}\|^{2}.$ (E.4)
Take $\eta_{1}=1-\kappa_{0}$ and $8\eta_{2}=\kappa_{0}-\kappa_{0}^{2}$, (E.4) becomes
$\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}\leq(2-2\sqrt{\kappa_{0}^{3}})\|\mathbf{x}\|^{2}.$ (E.5)
Therefore, for any $0<\delta_{0}<1$, take $\gamma$ as
@@
$1-\frac{\gamma^{2}}{4}=\sqrt[3]{\left(\frac{2-\delta_{0}^{2}}{2}\right)^{2}},$
and the result (E.5) becomes
$\text{dist}(\mathbf{x}_{{\mathcal{T}}},z)\leq\delta_{0}\|\mathbf{x}\|.$ (E.6)
Finally, we have
$\text{dist}(\mathbf{x},\mathbf{z})^{2}$~ $\makebox[29.6192pt][c]{$=$}~\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}+\|\mathbf{x}^{\perp}_{{\mathcal{T}}}\|^{2}$
$\makebox[29.6192pt][c]{$\overset{(a)}{\leq}$}~\delta_{0}^{2}\|\mathbf{x}\|^{2}+\frac{\gamma^{2}}{4}\|\mathbf{x}\|^{2}=\left(\delta_{0}^{2}+\frac{\gamma^{2}}{4}\right)\|\mathbf{x}\|^{2},$
where $(a)$ comes from (E.6) and the definition of $\mathcal{T}$}. Moreover, condition (51) indicates that $\delta\geq\sqrt{\delta_{0}^{2}+\frac{\gamma^{2}}{4}}\in[0,1]$. Therefore, we complete the proof of Proposition 4.

### v3

@@ Appendix B Proof for Proposition 1
Denote $\overline{S}$ as the index set corresponding to the largest $p$ elements of $\mathbf{x}$. We establish the upper and lower bounds for $\frac{1}{p}\sum_{j\in S^{0}}Z_{j}$ and $\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}$, respectively. Applying Lemma 6 yields
$\frac{1}{p}\sum_{j\in S^{0}}Z_{j}\leq$ $~\frac{1}{p}\sum_{j\in S^{0}}\mathbb{E}\left[\tilde{Z}_{j}\right]+\left\lvert\frac{1}{p}\sum_{j\in S^{0}}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
$\leq$ $~\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{4p\|\mathbf{x}\|^{2}}+\frac{1}{16ps(p)}$ (B.1)
and
$\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}\geq$ $~\frac{1}{p}\sum_{j\in\overline{S}}\mathbb{E}\left[\tilde{Z}_{j}\right]-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
$=$ $\frac{\|\mathbf{x}_{\overline{S}}\|^{2}}{4p\|\mathbf{x}\|^{2}}-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
@@
This completes the proof.
@@ Appendix C Proof for Proposition 2
For any set $\mathcal{T}$} satisfying $|{\mathcal{T}}|=p^{\prime}\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$, let $\mathbf{z}$ be the unit eigenvector corresponding to the largest eigenvalue, denoted as $\tau$, of
@@
where $(a)$ is based on the relationship between matrix norm and vector norm,
and $(b)$ employs Corollary 4. The above inequality holds with probability at least $1-n^{-c_{\eta}}$ when $m\geq C_{\eta}p^{\prime}s^{2}(p)\log n$ since $|{\mathcal{T}}|=p^{\prime}$. Therefore, through simple transformation, (C.6) can be expressed as
@@
Moreover, since $\tau$ is the largest eigenvalue of the Hermitian matrix $\mathbf{Z}_{{\mathcal{T}}}$, we can estimate it as
@@
where $(a)$ comes from the definition of matrix norm and Corollary 4. Taking (C.16) into (C.7) yields
@@
where $(a)$ comes from the assumption $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$. Let $\sqrt{1-16\eta}=1/2$, we complete the proof.
@@ Appendix D Proof for Proposition 3
$\overset{(a)}{\geq}$ $~\frac{1}{4\|\mathbf{x}\|^{2}}\times\frac{1}{2}\sqrt{\frac{1}{2s(p)}}\|\mathbf{x}\|\times\frac{\gamma\|\mathbf{x}\|}{2\sqrt{k}}-|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$ (D.7)
$\overset{(b)}{>}$ $~\frac{\gamma}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}}-\frac{\gamma}{32\sqrt{2}}\sqrt{\frac{1}{ks(p)}}=\frac{\gamma}{32\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.10)
where $(a)$ is from the definition of $S_{\gamma}$ and Proposition 2, and $(b)$ employs Lemma 7 with $\eta=\frac{1}{16\sqrt{2}}$ and holds with probability exceeding $1-n^{-c}$ when $m\geq C\gamma^{-2}ks(p)\log n$.
@@
$|f_{l}|\leq|\mathbb{E}[\tilde{f}_{l}]|+|f_{l}+\mathbb{E}[\tilde{f}_{l}]|\overset{(a)}{\leq}\frac{\gamma}{32\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.11)
where $(a)$ comes from (D.1) and Lemma 7 with $\eta=\frac{1}{16\sqrt{2}}$. By combining (D.10) and (D.11), we claim that all indices in $S_{\gamma}$ are selected in this case. Therefore, we can derive that
$\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\sum_{i\notin S_{\gamma}}\frac{|x_{i}|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}.$ (D.12)
@@ Appendix E Proof of Proposition 4
Recall that in Step 4 of the algorithm, $\mathbf{z}$ is defined as the eigenvector of $\mathbf{Z}_{S^{1}}$ corresponding to the largest eigenvalue, scaled such that $\|\mathbf{z}\|_{2}=\lambda$, where $\lambda=\sqrt{\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}}$. Let $\hat{\mathbf{u}}=\mathbf{z}/\|\mathbf{z}\|_{2}$ be the estimated direction, and $\mathbf{u}^{*}=\mathbf{x}_{S^{1}}/\|\mathbf{x}_{S^{1}}\|_{2}$ be the true direction on the subset $S^{1}$.
By the definition of the distance, we have
$\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})$ $=\min_{\phi\in[0,2\pi)}\|\mathbf{z}-e^{j\phi}\mathbf{x}_{S^{1}}\|_{2}$
$\leq\|\mathbf{z}-e^{j\phi^{*}}\mathbf{x}_{S^{1}}\|_{2}$
$=\Big\|\lambda\hat{\mathbf{u}}-e^{j\phi^{*}}\|\mathbf{x}_{S^{1}}\|_{2}\mathbf{u}^{*}\Big\|_{2},$ (E.1)
where $\phi^{*}$ is the optimal phase aligning $\hat{\mathbf{u}}$ and $\mathbf{u}^{*}$. By adding and subtracting $\lambda e^{j\phi^{*}}\mathbf{u}^{*}$, we can decompose the error into a norm component and a directional component via the triangle inequality:
$~~~~\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})$
$\leq\Big\|\lambda\hat{\mathbf{u}}-\lambda e^{j\phi^{*}}\mathbf{u}^{*}\Big\|_{2}+\Big\|\lambda e^{j\phi^{*}}\mathbf{u}^{*}-\|\mathbf{x}_{S^{1}}\|_{2}e^{j\phi^{*}}\mathbf{u}^{*}\Big\|_{2}$
$=\lambda\underbrace{\|\hat{\mathbf{u}}-e^{j\phi^{*}}\mathbf{u}^{*}\|_{2}}_{\text{directional error}}+\underbrace{|\lambda-\|\mathbf{x}_{S^{1}}\|_{2}|}_{\text{norm error}}.$ (E.2)
On one hand, for the directional error, the expected matrix is $\mathbb{E}[\tilde{\mathbf{Z}}_{S^{1}}]=\frac{\mathbf{x}_{S^{1}}\mathbf{x}_{S^{1}}^{*}}{4\|\mathbf{x}\|^{2}}$. The eigengap between the largest eigenvalue and the second largest (which is 0) is $\delta_{gap}=\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{4\|\mathbf{x}\|^{2}}$. According to the Davis-Kahan $\sin\Theta$ theorem [7], the distance between the eigenvectors is bounded by the ratio of the perturbation to the eigengap:
$\|\hat{\mathbf{u}}-e^{j\phi^{*}}\mathbf{u}^{*}\|_{2}$ $\leq\frac{\sqrt{2}\|\mathbf{Z}_{S^{1}}-\mathbb{E}[\tilde{\mathbf{Z}}_{S^{1}}]\|}{\|\mathbf{x}_{S^{1}}\|^{2}/(4\|\mathbf{x}\|^{2})}$
$\overset{(a)}{\leq}\frac{4\sqrt{2}\|\mathbf{x}\|^{2}}{\|\mathbf{x}_{S^{1}}\|^{2}}C\sqrt{\frac{k\log n}{m}},$ (E.3)
where $(a)$ comes from Corollary 4, with probability at least $1-2e^{-ck\log n}$.
On the other hand, for the norm error, we need to bound $|\lambda-\|\mathbf{x}_{S^{1}}\|_{2}|$. By the triangle inequality, we have
$|\lambda-\|\mathbf{x}_{S^{1}}\|_{2}|\leq|\lambda-\|\mathbf{x}\|_{2}|+|\|\mathbf{x}\|_{2}-\|\mathbf{x}_{S^{1}}\|_{2}|.$ (E.4)
First, from Lemma 2, we know that $|\lambda^{2}-\|\mathbf{x}\|^{2}|\leq C\sqrt{\frac{\log n}{m}}\|\mathbf{x}\|^{2}$. Using the inequality $|\sqrt{a}-\sqrt{b}|\leq\frac{|a-b|}{\sqrt{a}}$, we have
$|\lambda-\|\mathbf{x}\|_{2}|\leq\frac{|\lambda^{2}-\|\mathbf{x}\|^{2}|}{\|\mathbf{x}\|_{2}}\leq C\sqrt{\frac{\log n}{m}}\|\mathbf{x}\|_{2}.$ (E.5)
When $m>c\log n$, it holds that
$\lambda\leq\left(1+C\sqrt{\frac{\log n}{m}}\right)\|\mathbf{x}\|_{2}\leq 2\|\mathbf{x}\|_{2}.$ (E.6)
Second, from Proposition 3, we know $S^{1}$ contains the dominant entries $S_{\gamma}$, implying
$\|\mathbf{x}_{S^{1}}\|_{2}\geq\sqrt{1-\gamma^{2}/4}\|\mathbf{x}\|_{2}.$ (E.7)
We can always choose small $\gamma\leq 1/2$ such that $\sqrt{1-\gamma^{2}/4}\geq 1-\gamma/2$. Thus, we have
$|\|\mathbf{x}\|_{2}-\|\mathbf{x}_{S^{1}}\|_{2}|\leq\frac{\gamma}{2}\|\mathbf{x}\|_{2}.$ (E.8)
Collecting the above bounds (E.3) – (E.8) with constants absorbed into $C_{1}$, $C_{2}$, we have
@@
$~~~~\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})$ $\leq C_{1}\left(1-\frac{\gamma^{2}}{4}\right)^{-1}\|\mathbf{x}\|_{2}\sqrt{\frac{k\log n}{m}}$
$+$~ $\left(C_{2}\sqrt{\frac{\log n}{m}}+\frac{\gamma}{2}\right)\|\mathbf{x}\|_{2}.$ (E.9)
Choose a small constant $\gamma$ such that $\left(1-\frac{\gamma^{2}}{4}\right)^{-1}\leq 2$. We have
$\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})\leq C\sqrt{\frac{k\log n}{m}}\|\mathbf{x}\|_{2}+\frac{\gamma}{2}\|\mathbf{x}\|_{2}.$ (E.10)
Finally, considering the total distance to the original signal $\mathbf{x}$:
$\text{dist}(\mathbf{z},\mathbf{x})$ $\leq\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})+\|\mathbf{x}_{S^{1}}-\mathbf{x}\|_{2}$
$\leq C\|\mathbf{x}\|\sqrt{\frac{k\log n}{m}}+\gamma\|\mathbf{x}\|.$ (E.11)
To ensure $\text{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$, we set $\gamma<\frac{1}{2}\delta$. Proposition 3 implies a scaling dependency of $m\propto\gamma^{-2}$. Consequently, setting $\gamma<\frac{1}{2}\delta$ requires $m\propto\delta^{-2}$. Crucially, this sample complexity is sufficient to bound the first term in (E.11) by $\frac{1}{2}\delta\|\mathbf{x}\|$ as well. Thus, the condition is satisfied provided that
$m\geq C\delta^{-2}k\log n.$ (E.12)